OpenStudy (esshotwired):

Medal for correct answers. 2+2= 7+7= 3+3= Hint: The answers are not what you think they are. meaning they are not 4, 14, and 6

OpenStudy (esshotwired):

fyi: this is for fun for a free medal

undeadknight26 (undeadknight26):

lol maybe 22 77 and 33?

undeadknight26 (undeadknight26):

just a guess :D

OpenStudy (esshotwired):

not exactly what i was going for, but if no one answers i will give u the medal

undeadknight26 (undeadknight26):

yay! Lol pm me da answer :D

OpenStudy (uri):

2+2=fish 7+7=triangle 3+3=eight.

OpenStudy (esshotwired):

yes!

OpenStudy (anonymous):

@uri, may i talk to you?

OpenStudy (uri):

suree @pandaluvs

undeadknight26 (undeadknight26):

cute panda pictures!!!!!! look so cuddly like a teddy bear :D